### v1

# A Fast and Accurate Approach for Covariance Matrix Construction

Abstract
Reichel (2025) defined the bariance as $\operatorname{Bariance}(x)=\tfrac{1}{n(n-1)}\sum_{i<j}(x_{i}-x_{j})^{2}$, which admits an $\mathcal{O}(n)$ reformulation using scalar sums. We extend this to the covariance matrix by showing that $\operatorname{Cov}(\mathbf{X})=\tfrac{1}{n-1}\!\left(\mathbf{X}^{\top}\mathbf{X}-\tfrac{1}{n}\mathbf{s}\,\mathbf{s}^{\top}\right)$ with $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$ is algebraically identical to the pairwise-difference form yet avoids explicit centering. Computation reduces to a single $p\times p$ outer matrix product and one subtraction. Empirical Benchmarks in Python show clear runtime gains over numpy.cov and in non-BLAS tuned settings. Faster Gram routines such as RXTX [16] and for $\mathbf{X}\mathbf{X}^{\top}$ further reduce total cost.
MSC (2020): Primary 62H12; Secondary 62-08, 65F30, 65Y20, 15A09
1 Introduction
Reichel (2025) introduced the bariance, a between-sample variance based on all pairwise differences and computable in $\mathcal{O}(n)$ using scalar sums only. The bariance—the average of all pairwise squared differences—admits an algebraically optimized $\mathcal{O}(n)$ formula that is numerically equivalent to the naïve $\mathcal{O}(n^{2})$ definition and, for data, equals exactly twice the unbiased sample variance.
This work extends the same principle to the covariance and variance–covariance matrices. Starting from the pairwise-difference representation, we derive an optimized closed form that depends only on the Gram matrix $\mathbf{X}^{\top}\mathbf{X}$, the column-sum vector $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$, and the sample size $n$. Complete elementwise and matrix proofs are given, including an identity through the centering matrix $\mathbf{H}=\mathbf{I}_{n}-\tfrac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$, together with symbolic verification of all scalar equalities.
On the computational side, the optimized form avoids explicit centering, reduces data movement, and concentrates work in a single $p\times p$ outer-product matrix followed by one subtraction. The resulting estimator remains algebraically identical to the textbook sample covariance with denominator n1n-1 while achieving lower asymptotic cost.
Benchmarks in Python and R—with per-estimator warm-up, Tukey’s rules (IQR) trimming, and bootstrapped confidence bands—show that the optimized construction is slightly faster than a standard “center-then-multiply” computation and clearly faster than numpy.cov for large $n$, while remaining numerically identical up to machine precision. Comparisons with cov in base R confirm similar behavior, though the gain may diminish in BLAS-accelerated settings.
We also reference faster Gram-matrix routines such as RXTX [16] and for $\mathbf{X}\mathbf{X}^{\top}$, which can serve as a drop-in improvement or computational investment. Overall, the proposed estimator performs better in non–BLAS-tuned environments while preserving full analytical equivalence. The paper closes with a discussion of uses in standard-error computation and applications or generalized least-squares (GLS) estimation.
2 Background and notation
Assume $n\geq 2$ and that all variables are real. Pairwise variance decompositions appear in finite-population work [5]. Matrix and scalar-sum identities follow [17, 9, 11, 13]. For cost and numerical stability, we follow [6, 10, 7].
Let $x_{1},\dots,x_{n}\in\mathbb{R}$ and $y_{1},\dots,y_{n}\in\mathbb{R}$. Define scalar sums:
$S_{x}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}x_{i},\quad S_{y}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}y_{i},\quad S_{xx}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}x_{i}^{2},\quad S_{yy}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}y_{i}^{2},\quad S_{xy}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}x_{i}y_{i}.$
Sample means:
$\bar{x}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{S_{x}}{n},\qquad\bar{y}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{S_{y}}{n}.$
3 Algebraic derivation of the between-variance: Bariance
For $n\geq 2$:
Definition 3.1 (Bariance defined as in [15]).
$\operatorname{Bar}(x_{1},\dots,x_{n})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{2n(n-1)}\sum_{i\neq j}(x_{i}-x_{j})^{2}.$
Lemma 3.2 (Double-sum expansion).
$\sum_{i\neq j}(x_{i}-x_{j})^{2}=\sum_{i\neq j}x_{i}^{2}-2\sum_{i\neq j}x_{i}x_{j}+\sum_{i\neq j}x_{j}^{2}.$
Proof.
Expand $(x_{i}-x_{j})^{2}=x_{i}^{2}-2x_{i}x_{j}+x_{j}^{2}$ and sum over $i\neq j$. ∎
Lemma 3.3 (Counting identities).
$\sum_{i\neq j}x_{i}^{2}$ $=(n-1)\sum_{i=1}^{n}x_{i}^{2}=(n-1)S_{xx},$
$\sum_{i\neq j}x_{j}^{2}$ $=(n-1)S_{xx},$
$\sum_{i\neq j}x_{i}x_{j}$ $=\left(\sum_{i=1}^{n}x_{i}\right)\left(\sum_{j=1}^{n}x_{j}\right)-\sum_{i=1}^{n}x_{i}^{2}=S_{x}^{2}-S_{xx}.$
Proof.
For the first line, fix $i$ and count n1n-1 terms. The second line is symmetric. For the third, remove diagonal terms from the full double sum. ∎
Proposition 3.4 (Optimized Bariance identity).
$\operatorname{Bar}(x)=\frac{nS_{xx}-S_{x}^{2}}{n(n-1)}.$
Proof.
Insert Lemma 3.3 into Lemma 3.2:
$\sum_{i\neq j}(x_{i}-x_{j})^{2}=(n-1)S_{xx}-2(S_{x}^{2}-S_{xx})+(n-1)S_{xx}=2nS_{xx}-2S_{x}^{2}.$
Divide by 2n(n1)2n(n-1). ∎
Corollary 3.5 (Relation to unbiased variance).
$\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}=\operatorname{Bar}(x).$
Proof.
Expand:
$\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}=\sum_{i}x_{i}^{2}-2\bar{x}\sum_{i}x_{i}+n\bar{x}^{2}=S_{xx}-\frac{S_{x}^{2}}{n}.$
Divide by n1n-1: $\frac{S_{xx}-S_{x}^{2}/n}{n-1}=\frac{nS_{xx}-S_{x}^{2}}{n(n-1)}$, equal to Proposition 3.4. ∎
Some Properties. $\operatorname{Bar}(x+c)=\operatorname{Bar}(x)$ for constant $c$; $\operatorname{Bar}(ax)=a^{2}\operatorname{Bar}(x)$ for scalar $a$.
4 Algebraic derivation of the between-covariance
For $n\geq 2$:
Definition 4.1 (Between-covariance).
$C_{xy}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{2n(n-1)}\sum_{i\neq j}(x_{i}-x_{j})(y_{i}-y_{j}).$
Lemma 4.2 (Four-term split).
$\sum_{i\neq j}(x_{i}-x_{j})(y_{i}-y_{j})=\sum_{i\neq j}x_{i}y_{i}-\sum_{i\neq j}x_{i}y_{j}-\sum_{i\neq j}x_{j}y_{i}+\sum_{i\neq j}x_{j}y_{j}.$
Proof.
Expand and sum. ∎
Lemma 4.3 (Counting for mixed terms).
$\sum_{i\neq j}x_{i}y_{i}$ $=(n-1)\sum_{i=1}^{n}x_{i}y_{i}=(n-1)S_{xy},$
$\sum_{i\neq j}x_{j}y_{j}$ $=(n-1)S_{xy},$
$\sum_{i\neq j}x_{i}y_{j}$ $=\left(\sum_{i}x_{i}\right)\left(\sum_{j}y_{j}\right)-\sum_{i}x_{i}y_{i}=S_{x}S_{y}-S_{xy},$
$\sum_{i\neq j}x_{j}y_{i}$ $=S_{x}S_{y}-S_{xy}.$
Proof.
Apply the same logic as Lemma 3.3. ∎
Proposition 4.4 (Optimized pairwise covariance).
$C_{xy}=\frac{nS_{xy}-S_{x}S_{y}}{n(n-1)}.$
Proof.
Insert Lemma 4.3 into Lemma 4.2:
$\sum_{i\neq j}(x_{i}-x_{j})(y_{i}-y_{j})=2nS_{xy}-2S_{x}S_{y}.$
Divide by 2n(n1)2n(n-1). ∎
Theorem 4.5 (Equivalence to the textbook covariance).
$\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})=\frac{nS_{xy}-S_{x}S_{y}}{n(n-1)}.$
Proof.
Expand:
$\sum_{i}(x_{i}-\bar{x})(y_{i}-\bar{y})=\sum_{i}x_{i}y_{i}-\bar{x}\sum_{i}y_{i}-\bar{y}\sum_{i}x_{i}+n\bar{x}\bar{y}=S_{xy}-\frac{S_{x}S_{y}}{n}.$
Divide by n1n-1: $\frac{S_{xy}-S_{x}S_{y}/n}{n-1}=\frac{nS_{xy}-S_{x}S_{y}}{n(n-1)}$. ∎
Remark 4.6 (Symbolic Sanity Check).
Define symbols $n,S_{x},S_{y},S_{xx},S_{xy}$. Both
$\frac{S_{xx}-S_{x}^{2}/n}{n-1}-\frac{nS_{xx}-S_{x}^{2}}{n(n-1)},\qquad\frac{S_{xy}-S_{x}S_{y}/n}{n-1}-\frac{nS_{xy}-S_{x}S_{y}}{n(n-1)}$
simplify to zero.
5 Matrix formulation: elementwise and compact form
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ with entries $x_{ik}$. Define
$\mathbf{s}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{X}^{\top}\mathbf{1}_{n}\in\mathbb{R}^{p},\qquad\mathbf{G}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{p\times p}.$
Entrywise,
$s_{k}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}x_{ik},\qquad G_{k\ell}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}x_{ik}x_{i\ell}.$
Theorem 5.1 (Matrix covariance identity).
The unbiased covariance of the columns of $\mathbf{X}$ equals
$\bm{\Sigma}(\mathbf{X})=\frac{1}{n(n-1)}\bigl(n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top}\bigr).$ (1)
Entrywise proof.
For columns $k,\ell$, Theorem 4.5 with $x_{i}\leftarrow x_{ik}$ and $y_{i}\leftarrow x_{i\ell}$ gives
$\Sigma_{k\ell}=\frac{nG_{k\ell}-s_{k}s_{\ell}}{n(n-1)}.$
Stacking these entries yields (4). ∎
Theorem 5.2 (Equivalence achieved via the centering matrix).
Let $\mathbf{H}=\mathbf{I}_{n}-\tfrac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. Then
$\frac{1}{n-1}(\mathbf{X}^{\top}\mathbf{H}\mathbf{X})=\frac{1}{n(n-1)}\bigl(n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top}\bigr).$
Proof.
Use $\mathbf{H}=\mathbf{I}_{n}-\tfrac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. Then
$\mathbf{X}^{\top}\mathbf{H}\mathbf{X}=\mathbf{X}^{\top}\mathbf{X}-\frac{1}{n}\mathbf{X}^{\top}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\mathbf{X}=\mathbf{X}^{\top}\mathbf{X}-\frac{1}{n}\mathbf{s}\mathbf{s}^{\top}.$
Multiply by $\tfrac{1}{n-1}$ to obtain $\tfrac{n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top}}{n(n-1)}$. ∎
Remark 5.3.
The two theorems confirm that the bariance-style matrix equals the centered covariance in both forms. $\bm{\Sigma}(\mathbf{X})$ is symmetric and positive semidefinite with $\mathrm{rank}(\bm{\Sigma})\leq\min(p,n-1)$.
6 Cost model and where time is saved
The centered formulation first computes the mean vector $\bar{\mathbf{x}}$, forms the centered matrix $\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}$ (reading and writing $O(np)$ doubles), and then multiplies the result. The bariance-style formulation computes $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$, the Gram matrix $\mathbf{G}=\mathbf{X}^{\top}\mathbf{X}$, and performs one $p\times p$ outer product and one subtraction:
$\bm{\Sigma}_{\text{Bar}}=\frac{1}{n(n-1)}\bigl(n\mathbf{G}-\mathbf{s}\mathbf{s}^{\top}\bigr).$
Memory traffic decreases because no intermediate $n\times p$ array is stored. The main computational cost lies in forming $\mathbf{X}^{\top}\mathbf{X}$, which can rely on tuned BLAS level-3 kernels [7, 6, 10]. When $\mathbf{X}\mathbf{X}^{\top}$ is required instead, RXTX [16] lowers the multiplication count by about five percent, even for moderate $n,p$.
7 Benchmark protocol
For each matrices with parameter pair $(n,p)$:
  1. 1.
    Generate data $\mathbf{X}\sim\mathcal{N}(0,1)^{n\times p}$.
  2. 2.
    For each estimator, perform one warm-up call to stabilize caches and JIT compilation.
  3. 3.
    Record runtimes over multiple repetitions.
  4. 4.
    Remove outliers using the $1.5\timesIQR$ rule for each method and size.
  5. 5.
    Form bootstrap percentile bands (95%).
  6. 6.
    Report mean runtimes after trimming.
The estimators compared are:
  • Centered form:
    $\bm{\Sigma}_{\text{Ctr}}=\frac{1}{n-1}(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top})^{\top}(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top});$
  • Bariance-style form:
    $\bm{\Sigma}_{\text{Bar}}=\frac{1}{n(n-1)}(n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top});$
  • Built-in: $\texttt{numpy.cov}(\mathbf{X},\text{rowvar=False},\text{ddof}=1)$.
All experiments were conducted in double precision, using consistent random seeds across methods.
8 Numerical equivalence in finite precision
For each simulated matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, compute
$\bm{\Sigma}_{\mathrm{Bar}}=\frac{1}{n(n-1)}\!\left(n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top}\right),\qquad\bm{\Sigma}_{\mathrm{Ctr}}=\frac{1}{n-1}\!\left(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}\right)^{\top}\!\left(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}\right),$
where $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$ and $\bar{\mathbf{x}}=\tfrac{1}{n}\mathbf{s}$. The entrywise deviation
$\Delta_{\max}=\|\bm{\Sigma}_{\mathrm{Bar}}-\bm{\Sigma}_{\mathrm{Ctr}}\|_{\max}$
was recorded for various $(n,p)$ combinations. Across all tested sizes, the maximum difference stayed below $10^{-12}$ in IEEE 754 double precision, consistent with rounding and cancellation limits reported in [10]. These outcomes confirm that the algebraic equality between the bariance-style and centered formulations holds to full double-precision accuracy.
9 Benchmark figures
Figure 1: Runtime vs. $n$ for $p=10$. Warm-up, IQR trimming, and 95% bootstrap bands.
Figure 2: Runtime vs. $n$ for $p=50$. Same protocol.
Figure 3: Runtime vs. $p$ at fixed $n=4000$.
Figure 4: Bariance-only runtime for tall ($p=10$) and wide ($p=200$) matrices across several $n$.
Figure 5: Relative speed up ratios for $p=10$: centered/bariance and NumPy/bariance. Values >1>\!1 favor the bariance form.
Figure 6: Numerical agreement: $\max_{k,\ell}|\Sigma_{\text{Bar}}-\Sigma_{\text{Ctr}}|$ for $p=10$ across many draws. Values stay at floating-point noise.
Figure 7: Fixed-effects score-matrix covariance timing, warmed and IQR-cleaned.
10 Numerical equivalence in finite precision
For each simulated matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, compute
$\bm{\Sigma}_{\mathrm{Bar}}=\frac{1}{n(n-1)}\!\left(n\mathbf{X}^{\top}\mathbf{X}-\mathbf{s}\mathbf{s}^{\top}\right),\qquad\bm{\Sigma}_{\mathrm{Ctr}}=\frac{1}{n-1}\!\left(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}\right)^{\top}\!\left(\mathbf{X}-\mathbf{1}_{n}\bar{\mathbf{x}}^{\top}\right),$
where $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$ and $\bar{\mathbf{x}}=\tfrac{1}{n}\mathbf{s}$. The entrywise deviation
$\Delta_{\max}=\|\bm{\Sigma}_{\mathrm{Bar}}-\bm{\Sigma}_{\mathrm{Ctr}}\|_{\max}$
was recorded for a range of $(n,p)$ combinations. Across all tested sizes, the maximum difference stayed below $10^{-12}$ in IEEE 754 double precision, consistent with rounding and cancellation limits reported in [10]. These findings confirm that the algebraic equality between the bariance-style and centered forms holds to full double-precision accuracy.
11 Applications
Sandwich covariances.
Let $\mathbf{g}_{i}\in\mathbb{R}^{p}$ denote the score vector for observation $i$, stacked as rows in $\mathbf{G}\in\mathbb{R}^{n\times p}$, and let $\mathbf{H}=\mathbf{I}_{n}-\tfrac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. The empirical covariance of the scores can be written as
$\widehat{\bm{\Omega}}_{\text{sandwich}}=\frac{1}{n(n-1)}\!\left(n\,\mathbf{G}^{\top}\mathbf{G}-(\mathbf{G}^{\top}\mathbf{1}_{n})(\mathbf{G}^{\top}\mathbf{1}_{n})^{\top}\right)=\frac{1}{n-1}\,\mathbf{G}^{\top}\mathbf{H}\mathbf{G}.$ (2)
The bariance-style and centered forms coincide algebraically. Expression (2) arises in variance formulas for M-estimators and their sandwich extensions [12, 25, 8, 14, 4]. The bariance-style version depends only on $\mathbf{G}^{\top}\mathbf{G}$ and $\mathbf{G}^{\top}\mathbf{1}_{n}$ and does not require forming $\mathbf{H}\mathbf{G}$.
Panel and fixed effects.
For each unit in a panel with $T$ periods, the within transformation is $\mathbf{M}=\mathbf{I}_{T}-\tfrac{1}{T}\mathbf{1}_{T}\mathbf{1}_{T}^{\top}$. Given $\mathbf{X}_{i}\in\mathbb{R}^{T\times p}$, the covariance after demeaning satisfies
$\widehat{\bm{\Sigma}}_{i}=\frac{1}{T-1}\,\mathbf{X}_{i}^{\top}\mathbf{M}\mathbf{X}_{i}=\frac{1}{T-1}\!\left(\mathbf{X}_{i}^{\top}\mathbf{X}_{i}-\frac{1}{T}\mathbf{s}_{i}\mathbf{s}_{i}^{\top}\right),\quad\mathbf{s}_{i}=\mathbf{X}_{i}^{\top}\mathbf{1}_{T}.$
This bariance-style identity applies to each block and matches the algebra of the within estimator [2, 26].
Just-in-Time (JIT) streaming.
For sequential data, maintain cumulative quantities
$\mathbf{S}_{t}=\sum_{i=1}^{t}\mathbf{x}_{i},\qquad\mathbf{G}_{t}=\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}.$
The unbiased covariance at time $t\geq 2$ is
$\bm{\Sigma}_{t}=\frac{1}{t(t-1)}\!\left(t\mathbf{G}_{t}-\mathbf{S}_{t}\mathbf{S}_{t}^{\top}\right)=\frac{1}{t-1}\!\left(\frac{1}{t}\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\top}-\bar{\mathbf{x}}_{t}\bar{\mathbf{x}}_{t}^{\top}\right),\quad\bar{\mathbf{x}}_{t}=\tfrac{1}{t}\mathbf{S}_{t}.$ (3)
Each update adds one outer product $\mathbf{x}_{t}\mathbf{x}_{t}^{\top}$ and one rank-one correction, with cost $\mathcal{O}(p^{2})$. Fast matrix routines for $\mathbf{X}\mathbf{X}^{\top}$ such as RXTX [16] provide measurable computational savings [1, 24].
Bootstrap and resampling.
For a resample with multiplicity weights $\mathbf{w}\in\mathbb{N}_{0}^{n}$ and $\mathbf{W}=\operatorname{diag}(\mathbf{w})$, the covariance of the resample is
$\widehat{\bm{\Sigma}}^{\ast}=\frac{1}{n^{\ast}-1}\!\left(\mathbf{X}^{\top}\mathbf{W}\mathbf{X}-\frac{1}{n^{\ast}}(\mathbf{X}^{\top}\mathbf{w})(\mathbf{X}^{\top}\mathbf{w})^{\top}\right),\quad n^{\ast}=\sum_{i=1}^{n}w_{i}.$
Each resample depends only on the weighted Gram pair $(\mathbf{X}^{\top}\mathbf{W}\mathbf{X},\,\mathbf{X}^{\top}\mathbf{w})$. This form is practical when data are stored in aggregated or Gram form and aligns with generalized-inverse frameworks [3].
Appendix A Elementwise derivations using matrix entries
For clarity, write out the $(k,\ell)$ entry of the covariance estimator from (4):
$\Sigma_{k\ell}=\frac{1}{n(n-1)}\left(n\sum_{i=1}^{n}x_{ik}x_{i\ell}-\Big(\sum_{i=1}^{n}x_{ik}\Big)\Big(\sum_{i=1}^{n}x_{i\ell}\Big)\right).$ (4)
The centered definition is
$\frac{1}{n-1}\sum_{i=1}^{n}(x_{ik}-\bar{x}_{k})(x_{i\ell}-\bar{x}_{\ell})=\frac{1}{n-1}\left(\sum_{i=1}^{n}x_{ik}x_{i\ell}-\bar{x}_{k}\sum_{i=1}^{n}x_{i\ell}-\bar{x}_{\ell}\sum_{i=1}^{n}x_{ik}+n\,\bar{x}_{k}\bar{x}_{\ell}\right),$ (5)
where $\bar{x}_{k}=(\sum_{i}x_{ik})/n$ and $\bar{x}_{\ell}=(\sum_{i}x_{i\ell})/n$.
Elementwise derivation.
Substitute $\bar{x}_{k}$ and $\bar{x}_{\ell}$ into (5):
$\frac{1}{n-1}\left(\sum_{i=1}^{n}x_{ik}x_{i\ell}-\frac{1}{n}\Big(\sum_{i=1}^{n}x_{ik}\Big)\Big(\sum_{i=1}^{n}x_{i\ell}\Big)-\frac{1}{n}\Big(\sum_{i=1}^{n}x_{i\ell}\Big)\Big(\sum_{i=1}^{n}x_{ik}\Big)+\frac{n}{n^{2}}\Big(\sum_{i=1}^{n}x_{ik}\Big)\Big(\sum_{i=1}^{n}x_{i\ell}\Big)\right).$
Combine and simplify:
$\frac{1}{n-1}\left(\sum_{i=1}^{n}x_{ik}x_{i\ell}-\frac{1}{n}\Big(\sum_{i=1}^{n}x_{ik}\Big)\Big(\sum_{i=1}^{n}x_{i\ell}\Big)\right)=\frac{1}{n(n-1)}\left(n\sum_{i=1}^{n}x_{ik}x_{i\ell}-\Big(\sum_{i=1}^{n}x_{ik}\Big)\Big(\sum_{i=1}^{n}x_{i\ell}\Big)\right),$
matching (4). This completes the elementwise equivalence.
Matrix form.
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{s}=\mathbf{X}^{\top}\mathbf{1}_{n}$, and $\mathbf{H}=\mathbf{I}_{n}-\tfrac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$. The covariance matrix is
$\widehat{\bm{\Sigma}}=\frac{1}{n-1}\,\mathbf{X}^{\top}\mathbf{H}\mathbf{X}=\frac{1}{n-1}\left(\mathbf{X}^{\top}\mathbf{X}-\frac{1}{n}\mathbf{s}\,\mathbf{s}^{\top}\right).$ (6)
The $(k,\ell)$ entry of (6) equals (4), so the matrix and scalar derivations coincide.
Pairwise difference identity.
Starting from
$\widehat{\Sigma}_{k\ell}=\frac{1}{n(n-1)}\sum_{1\leq i<j\leq n}(x_{ik}-x_{jk})(x_{i\ell}-x_{j\ell}),$
use
$\sum_{i<j}a_{i}a_{j}=\tfrac{1}{2}\!\left[\!\left(\sum_{i}a_{i}\right)^{2}-\sum_{i}a_{i}^{2}\!\right],$
on columns $k$ and $\ell$. After expansion and cancellation one obtains (4), and from (6) the compact form
$\widehat{\bm{\Sigma}}=\frac{1}{n-1}\left(\mathbf{X}^{\top}\mathbf{X}-\frac{1}{n}\mathbf{s}\,\mathbf{s}^{\top}\right).$
Computational note.
Equation (6) requires one Gram product $\mathbf{X}^{\top}\mathbf{X}$ and one outer product $\mathbf{s}\mathbf{s}^{\top}$. This avoids constructing the centered matrix $\mathbf{H}\mathbf{X}$ and keeps the cost to a single $p\times p$ multiplication and one subtraction. Fast Gram routines such as RXTX [16] for $\mathbf{X}\mathbf{X}^{\top}$ can further reduce runtime in high-dimensional cases.
Appendix B Python and R references for replication
Python.
A minimal implementation (MVP) of the optimized covariance estimator is:
cov_bar(X) = (n * (X.T @ X) - np.outer(s, s)) / (n * (n - 1))
with $s=X^{\top}\mathbf{1}_{n}$ and $n=X.\texttt{shape[0]}$. This path triggers one level-3 BLAS multiply for $X^{\top}X$ (GEMM/DSYRK) and one rank-1 update for $s\,s^{\top}$ (GER), avoiding materialization of $X-\mathbf{1}_{n}\bar{x}^{\top}$. In many Python builds NumPy is linked to a generic or single-thread OpenBLAS; users often observe that raw @ calls are fast when BLAS is available, while higher-level helpers incur extra allocations and passes [19, 20]. Identifying the active BLAS and confirming that np.dot/@ dispatch into it is standard practice [19]. Under these conditions the closed form can outpace numpy.cov, which centers then multiplies and moves an $n\times p$ temporary.
R (base).
An (at least numerically) equivalent base-R implementation is:
bariance_cov <- function(X) {
  n <- nrow(X)
  s <- colSums(X)
  G <- crossprod(X)
  (n * G - tcrossprod(s, s)) / (n * (n - 1))
In R, crossprod/tcrossprod map directly to level-3 BLAS (e.g., DSYRK/DGEMM) and are already the preferred fast path [21, 22]. Guidance on BLAS-backed functions in base R points the same way [23]. Base cov() is a compiled routine that performs centering and Gram products through BLAS, so its inner loops remain inside C/Fortran once data pointers are set. With a multithreaded vendor BLAS, this often beats an R-level function wrapper, even if that wrapper calls crossprod internally.
Outcome.
Python: the closed form is faster when numpy.cov does extra allocations and the build is not strongly tuned; the single GEMM++GER path wins [19, 20]. R: cov() already leverages BLAS through compiled code; cov() remains faster than an R-level wrapper that still returns to the interpreter between BLAS calls [21, 22, 23].